OpenStudy (anonymous):

Find the constants m and b in the linear function. f(0) = 5 and f(3) = -4.

OpenStudy (anonymous):

we have two points (0,5) (3,-4) 

OpenStudy (anonymous):

so y=-3x+c to find c we can use any point lets use (0,5)

OpenStudy (anonymous):

y=-3x+c 5=-3(0)+c c=5 y=-3x+5 is ur equation of the line

OpenStudy (anonymous):

in this case c=b=5